### v1

# IrokoBench: A New Benchmark for African Languages in the Age of Large Language Models

Abstract
Despite the widespread adoption of Large language models (LLMs), their remarkable capabilities remain limited to a few high-resource languages. Additionally, many low-resource languages (e.g., African languages) are often evaluated only on basic text classification tasks due to the lack of appropriate or comprehensive benchmarks outside of high-resource languages. In this paper, we introduce IrokoBench—a human-translated benchmark dataset for 16 typologically-diverse low-resource African languages covering three tasks: natural language inference (AfriXNLI), mathematical reasoning (AfriMGSM), and multi-choice knowledge-based QA (AfriMMLU). We use IrokoBench to evaluate zero-shot, few-shot, and translate-test settings (where test sets are translated into English) across 10 open and four proprietary LLMs. Our evaluation reveals a significant performance gap between high-resource languages (such as English and French) and low-resource African languages. We observe a significant performance gap between open and proprietary models, with the highest performing open model, Aya-101 only at 58% of the best-performing proprietary model GPT-4o performance. Machine translating the test set to English before evaluation helped to close the gap for larger models that are English-centric, like LLaMa 3 70B. These findings suggest that more efforts are needed to develop and adapt LLMs for African languages.
1 Introduction
In recent years, the capabilities of large language models (LLMs) have greatly improved, from coherent chat experiences to solving complex and knowledge-intensive tasks like mathematical reasoning, coding, and question answering [47, 33, 26]. These models have also demonstrated the ability to quickly learn new and challenging tasks with few in-context learning examples and through chain-of-thought reasoning [17, 53, 64]. However, most state-of-the-art LLMs are primarily trained on high-resource languages, resulting in sub-optimal performance for languages unseen during pre-training [58, 45]. Furthermore, this language coverage bias is reflected in the evaluation stage, predominantly conducted in English and a few other high-resource languages.
There has been considerable effort to create benchmarks for African languages, but they typically cover simpler tasks, or are specific to narrow domains such as machine translation, and, more recently, reading comprehension [13, 11]. Some diverse reasoning benchmarks have included Swahili—the most spoken native African language—in benchmarks like commonsense reasoning [48] and natural language inference [21]. However, performance in Swahili does not reflect the performance on the over 2,000+ languages in Africa, which often lack substantial digital textual resources comparable to that available for Swahili. Consequently, current multilingual evaluations of LLMs do not accurately reflect capabilities in reasoning and knowledge-intensive tasks across the majority of African languages. Furthermore, the few comprehensive evaluations that exist across languages often rely on machine translation of English benchmarks [55]. While automatic translation is a popular approach given the cost and time investment required for human translation, it often suffers from noise and biases. For example, translated datasets often introduce artefacts from the translation model  [61, 28, 51] or fail to reflect cultural context appropriately [63, 32, 49]. Automatic curation may also amplify any of the ubiquitous issues with the quality of broad pretraining sets [39, 35, 24, 18].
In this paper, we seek to address both the diversity and breadth of evaluation coverage. We introduce IrokoBench, a human curated benchmark dataset for 16 typologically diverse African languages which encompasses three complex tasks: natural language inference (NLI), mathematical reasoning, and multi-choice knowledge-based QA. The datasets were created by human translating a subset of English cross-lingual NLI (XNLI) [21], English Multilingual Grade School Math (MGSM) [54], and Massive Multitask Language Understanding (MMLU) [29], evaluation datasets into each of the 16 languages using professional translators.
We conduct a large-scale evalution of IrokoBench to evaluate zero-shot, few-shot, and translate-test settings (where test sets are translated into English) across 10 open and four proprietary LLMs. Our main contributions can be enumerated as follows:
  1. 1.
    We introduce and release IrokoBench, a human-translated benchmark that includes languages from various geographical regions: six from West Africa, five from East Africa, four from Southern Africa, and one from Central Africa, all with varying degrees of “low-resourcedness.” [34].
  2. 2.
    Sharp cliff in performance across all models on low-resource languages Our evaluation shows a large gap ( $\sim$ $45$% on average) between the performance of high-resource languages (e.g., English) and African languages on all LLMs evaluated. Notably, Swahili performs better than other African languages, likely due to its frequent use as a benchmark.
  3. 3.
    Models generally perform poorly in in-language evaluation When evaluating models with prompts in native languages, state-of-the-art open models such as Gemma, Llama 3 are outperformed by massively multilingual models like Aya-101. This can be attributed to a lack of trained models to respond in the native languages of the users. Machine translating the test set to English before evaluation helped to close the gap for English-centric models; however, requiring users to always translate their prompts from their native language to English may not be desirable behavior.
  4. 4.
    IrokoBench highlights the performance divide between open and proprietary models on low-resource languages. We find that proprietary closed models generally outperform open models for African languages. However, even these proprietary models exhibit substantial performance drops, due to the limited monolingual web data for African languages. The lowest performance is observed in languages such as Ewe, Lingala, Luganda, Twi and Wolof, which each have less than 50 million characters of available data [36]. Among the tasks evaluated, AfriMGSM proves most challenging for LLMs, followed by AfriMMLU and AfriXNLI.
These results underline the need for focused development and adaptation of LLMs to better support African languages, especially those with limited data resources. We release IrokoBench on HuggingFace under the CC BY-SA 4.0 licence https://huggingface.co/collections/masakhane/irokobench-665a21b6d4714ed3f81af3b1 to further multilingual evaluation and research.
2 Related Work
Multilingual Evaluation of LLMs:
The evaluation of multilingual capabilities of LLMs has garnered significant attention. This has led to an increase in research that explores their performance across diverse linguistic landscapes [7, 8, 37, 31, 14, 60, 55]. Despite this growing interest, there remains a notable lack of representation of African low-resource languages in these studies. [45, 12] address a broader spectrum of African languages, aligning more closely with our research. However, their study focuses on conventional NLP tasks, such as text classification, named entity recognition, question answering, and text generation. To address the lack of difficult benchmarks, a few works automatically translated MMLU benchmarks [47, 38, 60], but they do propagate errors of machine translation (MT) engines. Moreover, this is not applicable to low-resource languages with low-quality MT systems [2, 57]. Our research advances this by evaluating LLMs on more complex tasks using newly developed, human-annotated benchmarks specifically for African languages.
African Benchmark Datasets:
Due to the limited representation of African languages in the field of NLP, there has been a growing effort to create benchmark datasets for African languages to enable research on these languages, especially evaluating existing NLP tools on these languages. Initiatives such as Masakhane have been instrumental in the creation of standard benchmark for tasks such as machine translation [2], named entity recognition [5, 3], part of speech tagging [23], news topic classification [6], and sentiment analysis [43]. There are also several multilingual benchmark datasets that cover a few African languages, such as SIB-200 [4], Flores [27, 57], Aya dataset and Collection [55] and Taxi1500 [40]. However, despite all these efforts, African languages still lack quality and more difficult datasets to evaluate the reasoning capabilities of LLMs, which have become ubiquitous. Hence, this work contributes to this space by creating datasets to evaluate general use LLMs in African languages in a variety of downstream tasks.
3 IrokoBench
Figure 1: Task Description for IrokoBench datasets. Both AfriMGSM and AfriMMLU focus on QA, while AfriXNLI focuses on natural language inference between two pairs of sentences. For clarity, this figure provides examples in English.
Table 1: The IrokoBench datasets: dataset name, number of African languages covered, data split, and the subjects or domains covered. We included English, French, and Swahili in all benchmarks.
No. of No. instances
Dataset languages Train/Dev/Test Subjects / Domains
AfriMGSM 15 (excl. Swahili) 8  /  -  /  250 grade school mathematics
AfriMMLU 17 (incl. French) 25  /  83  /  500 elementary mathematics, high-school geography, International law, global facts, high school microeconomics
AfriXNLI 15 (excl. Swahili) -  /  450  /  600 face-to-face, telephone, oxford university press (oup), fiction, travel, government, nineeleven, letters, slate, verbatim
3.1 Languages covered by IrokoBench
We cover 18 languages in our evaluation, 16 native African languages, and two European languages (English and French) which are widely spoken and official in many African countries. English and French are also the source languages we translated from. We cover 16 diverse and widely spoken African languages from four regions of Africa: six from West Africa (Ewe, Hausa, Igbo, Twi, Wolof, Yoruba), five from East Africa (Amharic, Kinyarwanda, Luganda, Swahili, and Oromo), four from Southern Africa (chiShona, isiXhosa, isiZulu, and Sesotho), and Central Africa (Lingala). The African languages are from two biggest language families in Africa: three from Afro-Asiatic and 13 from the Niger-Congo family—where we cover eight Bantu languages. Table 7 provides an overview of the languages covered, including their family, regions, and the number of native speakers.
3.2 Tasks covered by IrokoBench
The selection of these tasks is primarily driven by their coverage across various domains and downstream tasks for diverse use cases. Additionally, they enable the evaluation of logical, abstract, and reasoning capabilities in LLMs, which is the hallmark of human intelligence [16, 30]. Figure 1 provides examples of the different tasks covered in our datasets. We provide their descriptions below:
AfriXNLI
The task of NLI involves the classification of a pair of sentences—a premise and a hypothesis as to entailment, neutral, or contradiction semantic relation. For example, the sentence “so I’m not really sure why” contradicts “I am certain as to the reason why” but has a neutral relation to “I don’t know why he transferred schools”. Here, we human translate the English portion of XNLI (a multilingual dataset comprising 15 languages, including Swahili) into the 15 African languages (excluding Swahili). While the original XNLI dataset has over 2,500/5,000 as DEV/TEST split, Each language in AfriXNLI has only 450 DEV instances and 600 test instances. We selected an equal number of instances from the 10 domains of XNLI. The task is evaluated using the accuracy metric since the dataset has balanced classes.
AfriMMLU
This is a multi-choice knowledge QA curated from freely available online sources by undergraduate and graduate students in the USA. The subjects cover simple general knowledge questions like “global fact” to highly-technical questions like “professional law” and “professional medicine”. MMLU are often grouped into STEM, Humanities, Social Sciences, and Others category. We focus on five subjects that we believe present less tendency to be culturally unbiased and that are not too difficult to translate since many of the subjects covered are only taught in African countries using English or French, making it extremely difficult to translate highly technical subjects, especially the STEM subjects. Table 1 shows the five subjects covered: two social science subjects (high-school geography and high-school microeconomics), one STEM subject (elementary mathematics), one humanities subject (international law), and one OTHER category (global facts). In total, we translated 608 question-answer pairs, with 500 instances in the test split, 100 questions per subject. The task is evaluated using the option prediction accuracy.
AfriMGSM
This is a QA task with questions obtained from grade school mathematical word problems created by human problem writers. AfriMGSM expands the original MGSM dataset [54], which contains 250 QA pairs and 11 languages (including Swahili), to 15 more languages. The dataset consists of 8 training examples for few-shot and chain-of-thought prompting and 250 as a test set. We evaluated this task using the Exact Match metric, which is popularly used for QA tasks.
3.3 Data collection process
Translation
We recruited language coordinators for each of the 16 African languages and French, and asked them to recruit professional translators to translate the sentences. The translation process took about two months, they started with XNLI, then MGSM and MMLU. Each translator received an appropriate remuneration for their work.  [^1] We recruited a logistic company in Kenya that managed all recruitment and payments—each country has different rates. Most of the translators translated from English except for Ewe, Lingala and Wolof translators that translated from French since they are from the Francophone region of Africa. Additionally, we translated the MMLU dataset to French by professional translators and from French to these three languages. Many of the Francophone translators understand French and English but are more fluent in French, so they could cross-check from English if the French sentences were not clear enough.
Quality control
Regarding quality control, language coordinators reviewed and corrected any poorly translated sentences. Translators received payment only after this phase to ensure the quality of translations. For additional checks, we computed COMET [50] quality estimation (QE) scores between the human translation and the original sentences based on AfriCOMET QE metric [62]. In general, the distribution of the scores (range between 0 and 1) reflect that most translated sentences are between $0.7$ and $1.0$ for about 13 language pairs except for Lingala, Twi, and Wolof where the average is around $0.5$. Further analysis shows that we cannot rely on these scores for those three languages since they are not covered in the pre-training of the original AfroXLMR encoder [9] used to build the AfriCOMET QE metric. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) [62]. We provide further analysis of the COMET scores in Appendix A.3.
3.4 LLMs used for evaluation
Open LLMs
We evaluate on three encoder-decoder open LLM: Flan-T5-XXL [19], mT0-XXL-MT [42], and Aya-101 [59] that have been instruction fine-tuned on T5 (FlanT5) and multilingual T5 pre-trained on 101 languages (mT0 and Aya-101) models. Furthermore, these models are also all designed to be massively multilingual and explicitly optimized to work outside of English. The instruction data were constructed from several supervised NLP tasks with diverse prompts in different languages. Aya-101 additionally includes some synthetic data generated from Cohere Command model [20], and a large amount of translated data from English to cover several languages [55]. The languages covered during instruction tuning differ for different models, mT0, FlanT5, and Aya covered 46, 60, and 101 languages respectively.
Additionally we evaluate of seven decoder-only open LLM models: BLOOMZ 7B [65], Gemma 7B [56], LLaMa 2 7B [58], LLaMa 3 (8B & 70B) [41], Command-(R & R+) [20]. These models’ weights are openly available under various licenses, ranging from fully permissive to non-commercial, research-only licenses. We evaluate the instruction-tuned variant of these models. Generally, with the exception of BLOOMZ, these models did not report including any African languages in their pre-training. BLOOMZ included 22 African languages, mainly from Niger-Congo language family.
Closed LLM
We limit our evaluation to only OpenAI GPT (3.5-0125, 4-Turbo-0125, 4o) [46], and Claude OPUS [10] models. Recent work has shown that proprietary models tend to exhibit better multilingual capabilities [8], although specifics regarding their pre-training and instruction fine-tuning processes are not disclosed.
3.5 Evaluation settings
Evaluation Set-up
We conduct two types of evaluations: in-language evaluation and translate-test evaluation, where test instances are automatically translated into English using a machine translation engine. For automatic translation, we use NLLB-200 (3.3B) [57]. In both in-language and translate-test setups, we perform cross-lingual transfer experiments from English and zero-shot evaluations by prompting LLMs. Few-shot evaluations are performed only for the two best models (one open and one closed model) in the in-language setting.
We use the EleutherAI LM Evaluation Harness (lm-eval) tool [15]—a popular evaluation tool that is helping to standardize LLM evaluation, especially for open models on HuggingFace Model Hub. For closed models, except for the AfriMGSM task where we still use lm-eval, we employ a verbalizer [25, 52] for prediction and evaluation. We provide more details in Appendix A.2.
Cross-lingual transfer experiments
We first conduct a study on cross-lingual transfer in a supervised learning setting by fine-tuning the English training data ($400K$ instances) from  Conneau et al. [21] and evaluating on the remaining languages. This experiment focuses solely on the NLI task due to the availability of training data for supervised learning. The evaluation employs several masked language models, including XLM-R [22], Serengeti [1], AfroXLMR-{base, large} [9], AfroXLMR-76L [4]. We report the result of the best model in the paper and others in Appendix A.5.
Zero- and few-shot evaluation
In a zero-shot setting, we use the prompts detailed in Table 8. For few-shot evaluations, we conduct a 5-shot assessment for both AfriMMLU and AfriXNLI, and an 8-shot assessment for AfriMGSM.
Table 2: Main results: Average performance of various LLMs on AfriMMLU, AfriXNLI and AfriMGSM. Except for AfriMGSM, which uses the Exact Match metric, others use the Accuracy.
AfriXNLI AfriMMLU AfriMGSM Ave. Ave. Ave. Ave.
in- translate in- translate in- translate in- translate English French
Model size lang. test lang. test lang. test lang. test lang. lang.
AfroXLMR-76L 559M 65.7 63.6 65.7 63.6 88.2 83.3
FlanT5-XXL 11B 36.7 53.2 25.4 35.1 1.1 8.2 21.1 32.2 44.7 35.4
mT0-XXL-MT 13B 51.0 49.9 28.2 28.9 2.6 2.6 27.3 27.1 34.4 33.3
Aya-101 13B 49.9 49.4 30.3 31.1 3.3 6.6 27.9 29.0 36.2 35.0
BLOOMZ 7B 7B 39.4 47.5 23.5 26.4 4.1 2.1 22.4 25.3 32.0 29.9
Gemma 7B 7B 36.4 42.0 29.5 38.3 4.6 8.2 23.5 29.5 41.2 37.6
LLaMa 2 7B 7B 33.4 42.3 24.7 29.9 1.7 5.9 19.9 26.0 34.8 32.3
LLaMa 3 8B 8B 35.8 44.1 27.5 34.4 2.9 27.0 22.0 35.2 47.0 39.5
LLaMa 3 70B 70B 34.2 40.9 34.1 47.4 8.2 45.3 25.5 44.5 70.7 57.7
Command R 35B 34.0 46.1 27.0 36.5 1.6 5.8 20.9 29.5 42.1 38.4
Command R+ 104B 37.9 45.4 26.1 35.7 2.3 9.2 22.1 30.1 47.8 39.3
Claude OPUS UNK 58.1 56.4 43.0 47.6 25.9 32.7 42.3 45.6 73.3 64.0
GPT-3.5 Turbo UNK 41.5 52.2 31.8 39.6 5.3 19.3 26.2 37.0 64.7 54.1
GPT-4-Turbo UNK 60.0 52.8 50.9 47.0 26.8 25.1 45.9 41.6 81.7 74.1
GPT-4o UNK 62.7 59.9 52.3 52.3 29.3 32.4 48.1 48.2 72.5 65.1
Average 28.2 34.3 51.6 45.4
4 Results
Large performance gaps between high-resource languages and African languages
Table 2 shows the result of zero-shot evaluation for various LLMs. On average, there is a significant performance gap between African languages and English (up to $45\%$) and French (up to $36\%$). The best model for African languages is GPT-4o, with an average performance of 48.1 across the evaluated tasks. However, GPT-4-Turbo outperforms GPT-4o for English and French, showing improvements of 9.2 and 9.0, respectively, across all tasks. It appears that GPT-4o’s enhancements for low-resource languages may negatively impact its performance on high-resource languages. Finally, as shown in Table 7, the languages with the lowest performance have the least monolingual data on the web.
Large performance gaps between closed and open weights models
Our results, as presented in Table 2, indicate that the closed models Claude OPUS, GPT-4-Turbo, and GPT-4o consistently outperform the open models on our IrokoBench benchmarks. The closed models achieve average performance scores ranging from $42.3$ to $48.1$ across all tasks. The performance gap between the best closed model (GPT-4o) and the best open model (Aya-101) is $20.2$. Notably, the largest performance differences are observed in the AfriMMLU and AfriMGSM tasks, where GPT-4o outperforms LLaMa 3 70B by $18.2$ and $21.1$, respectively. For the AfriXNLI task, mT0-XXL-MT and Aya-101 perform better than LLaMa 3 70B and also achieve the highest scores among all open-weight models.
Majority models perform worse for in-language prompting Most users would prefer to prompt in their native language; however, we find that almost all models we benchmark perform better with prompts translated into English. Only a few exceptions, such as GPT-4-Turbo, mT0-XXL-MT, and Aya-101, perform better in the in-language evaluation. Among open-source models, Aya-101 performs the best in the in-language evaluation, surpassing much larger open-weight models like LLaMa 3 70B. Specifically, LLaMa 3 70B and LLaMa 3 8B benefit the most from the translate-test approach, showing average improvements of $19.0$ and $13.1$, respectively. Notably, LLaMa 3 70B achieves the best overall results for the AfriMGSM task with the translate-test, outperforming GPT-4o by more than $12.9$. Similarly, we observe significant performance gains for the FlanT5-XXL and GPT-3.5 Turbo models on the AfriXNLI and AfriMGSM tasks when using the translate-test. We attribute this boost to the fact that these models are heavily English centric.
Table 3: AfriMMLU results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
Aya 41.8 36.4 32.8 26.8 32.2 35.8 30.0 26.0 30.8 30.6 27.8 34.2 35.0 25.8 24.6 31.2 29.8 31.8 30.3
LLaMa 3 8B 49.0 45.6 26.8 25.8 29.6 29.4 28.4 28.8 22.4 29.2 29.4 27.0 30.4 24.8 27.6 24.6 28 28.0 27.5
LLaMa 3 70B 74.4 68.6 33.0 31.8 33.6 37.8 32.4 35.0 30.8 33.0 33.4 35.2 44.8 33.7 32.0 30.0 33.8 35.0 34.1
GPT-4 Turbo 80.7 76.5 37.7 23.9 55.2 54.5 54.9 46.8 40.9 39.8 54.1 55.4 61.0 30.1 31.0 56.2 48.9 58.4 50.9
GPT-4o 78.6 71.4 55.2 31.6 57.4 55.8 56.8 55.4 45.0 48.8 58.4 62.0 67.4 37.6 32.6 60.8 50.8 60.6 52.3
Translate-Test (Eval. in English)
Aya 37.2 32.0 30.4 30.8 31.6 29.2 32.6 29.2 29.0 33.4 33.8 31.6 30.8 25.4 32.8 32.6 32.4 31.1
LLaMa 3 8B 43.4 38.4 35.2 34.4 33.2 36.0 36.0 27.8 34.2 31.4 38.8 43.6 29.0 29.6 36.6 33.2 33.6 34.4
LLaMa 3 70B 64.4 54.4 42.2 44.4 48.6 47.2 49.8 41.2 49.6 48.4 54.0 60.0 41.4 29.8 49.6 49.2 49.2 47.4
GPT-4 Turbo 61.0 50.0 37.6 43.2 47.6 48.2 46.8 40.4 44.6 46.2 53.0 72.7 36.4 28.0 48.4 46.6 49.0 47.0
GPT-4o 67.2 57.6 40.6 49.4 56.4 55.0 50.2 41.8 51.6 49.8 59.0 80.0 37.2 33.0 53.0 52.4 54.2 52.3
Table 4: AfriMGSM results in in-language and translate-test scenarios: Exact Match score per language. Average computed on only African languages
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
Aya 24.8 22.0 7.6 2.8 5.6 4.0 6.8 2.0 4.4 2.4 3.2 3.2 11.2 1.2 2.4 2.0 3.6 4.0 4.2
LLaMa 3 8B 24.8 22.0 7.6 2.8 4.8 4.4 5.6 2.0 4.8 2.4 2.4 3.2 11.2 1.2 2.4 1.6 3.2 3.6 4.0
LLaMa 3 70B 86.4 54.4 5.2 2.8 15.6 9.2 8.4 5.6 8.0 4.0 4.8 6.0 34.8 3.2 4.0 4.4 7.2 7.6 8.2
GPT-4 Turbo 79.6 64.0 8.4 6.8 40.8 21.6 41.6 30.8 24.4 31.2 38.4 38.0 54.4 6.8 9.2 31.2 29.2 31.2 27.8
GPT-4o 58.4 43.6 4.4 5.6 42.0 32.0 36.4 30.8 22.8 39.2 33.2 28.8 44.8 10.4 12.4 29.2 30.8 28.4 27.0
Translate-Test (Eval. in English)
Aya 8.0 8.0 6.0 7.6 4.0 7.2 5.2 4.4 5.2 8.4 8.0 10.8 4.4 5.6 6.0 8.4 6.4 6.6
LLaMa 3 8B 44.0 31.6 18.0 27.2 24.0 34.0 27.2 23.2 28.8 26.8 33.6 38.0 19.6 12.8 28.8 29.6 28.8 27.0
LLaMa 3 70B 69.6 53.6 27.6 48.4 40.4 54.0 47.6 34.0 47.6 46.8 54.0 69.6 29.2 20.8 50.0 49.2 51.2 45.3
GPT-4 Turbo 63.6 42.0 20.4 39.2 35.2 47.6 41.6 27.2 38.8 39.2 46.0 56.8 23.2 12.4 41.2 42.0 40.4 37.1
GPT-4o 52.0 18.4 37.6 36.4 28.0 34.8 38.4 28.4 32.8 32.4 36.8 45.2 17.6 12.4 31.6 36.0 38.0 31.6
Table 5: AfriXNLI results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
AfroXLMR-76L 88.2 83.3 78.5 58.3 73.3 70.0 65.8 33.3 68.0 69.3 70.8 70.8 73.3 59.5 51.8 73.0 63.2 72.5 65.7
Aya 61.5 60.5 58.2 42.2 56.8 54.7 48.7 34.5 44.5 51.7 55.8 54.7 55.2 48.2 37.3 53.7 48.3 54.3 49.9
LLaMa 3 8B 54.0 51.8 34.5 33.8 36.0 39.5 33.8 32.5 35.5 34.7 37.8 35.0 39.5 33.0 33.5 38.1 38.1 36.8 35.8
LLaMa 3 70B 51.3 50.2 36.8 33.3 35.3 36.5 34.5 33.0 33.5 33.3 33.8 34.3 36.5 33.0 33.5 34.0 33.0 33.0 34.2
GPT-4 Turbo 86.0 82.2 65.2 38.7 69.3 69.5 67.7 32.2 65.0 58.3 67.5 65.7 67.2 47.5 39.7 72.0 65.3 69.2 60.0
GPT-4o 86.2 78.7 66.7 48.3 69.2 68.2 66.8 31.2 67.2 66.2 69.8 68.3 72.5 53.2 49.5 72.5 63.5 70.0 62.7
Translate-Test (Eval. in English)
AfroXLMR-76L 83.0 73.7 54.3 67.2 66.0 63.0 32.8 65.7 65.8 71.2 70.2 73.0 56.8 47.5 74.2 63.7 72.0 63.6
Aya 60.3 55.2 47.0 51.8 52.2 50.3 33.0 49.0 49.0 52.0 53.3 54.0 45.5 41.5 54.5 48.8 54.0 49.4
LLaMa 3 8B 52.2 48.3 43.3 43.2 44.8 45.3 33.7 46.0 45.2 45.5 47.5 47.5 40.5 38.7 44.5 44.2 47.0 44.1
LLaMa 3 70B 48.5 42.8 41.3 43.7 39.8 40.3 33.3 43.3 39.7 41.1 43.0 44.3 39.7 38.7 42.3 38.2 42.7 40.9
GPT-4 Turbo 75.2 64.0 43.0 52.8 59.0 55.8 31.7 51.0 52.5 55.8 58.5 60.2 45.0 39.7 63.2 52.5 59.5 52.8
GPT-4o 80.0 69.0 50.7 62.2 63.3 61.0 33.0 63.2 58.7 65.8 67.0 69.5 51.3 46.2 69.7 60.5 67.3 59.9
Table 6: Ablation results: Effect of using different prompts for AfriXNLI. Best results per prompt are in bold. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
ANLI prompt [44]—simple prompt
Aya 61.5 60.5 58.2 42.2 56.8 54.7 48.7 34.5 44.5 51.7 55.8 54.7 55.2 48.2 37.3 53.7 48.3 54.3 49.9
Gemma 7B 51.5 47.8 34.3 35.3 38.8 39 35.8 33.2 34.8 35.7 38.8 35.7 39.3 37.2 35.8 37.2 36.5 35.7 36.4
LLaMA 3 70B 51.3 50.2 36.8 33.3 35.3 36.5 34.5 33 33.5 33.3 33.8 34.3 36.5 33 33.5 34 33 33 34.2
Opus 69.3 67.5 54.8 41.3 55.5 52.8 46.8 27.8 51.3 52.7 50.2 56.8 57.8 46.7 41.0 57.7 53.3 54.5 50.1
GPT-4o 79.8 74.5 61.8 42.0 61.0 58.2 55.8 33.2 58.0 58.3 62.2 62.0 65.5 49.2 46.5 62.5 55.2 59.8 55.7
Lai’s prompt [37]—detailed prompt
Aya 49.4 45.1 47.1 32.3 44.2 43.8 39.1 21.5 40.3 38.5 41.7 43.7 43.7 38 27.3 46.2 39.8 43.2 39.4
Gemma 7B 64 58.2 19.5 19.1 24.2 25.8 22.2 24.2 20.2 20.6 25.8 20.3 46.7 19.6 19 31.1 20.3 25.8 24.0
LLaMA 3 70B 72.8 51.4 41.9 27.5 47.1 43 39.8 23.9 38 38.2 37.9 37.4 52.1 22.1 29.1 39 36.7 36.3 36.9
Opus 85.7 74.7 61.3 54.5 61.5 45.0 64.5 30.5 63.7 50.2 57.0 68.3 70.5 56.0 50.0 68.8 63.7 63.7 58.1
GPT-4o 86.2 78.7 66.7 48.3 69.2 68.2 66.8 31.2 67.2 66.2 69.8 68.3 72.5 53.2 49.5 72.5 63.5 70.0 62.7
4.1 Task-specific results
Here, we examine the individual language performance per task, comparing which task prefers in-language v.s. translate-test evaluation. We compare the performance on a subset of LLMs: Aya-101, LLaMa 3 (8B & 70B), GPT-4-Turbo, and GPT-4o. Other LLMs are in Appendix A.4.
AfriMMLU evaluation is better when prompting in-language
Table 3 shows the result of different LLMs on AfriMMLU using in-language vs. translate-test. We found in-language prompting to be better on 10 out of the 16 African languages we evaluated; ibo, orm, wol, and yor have comparable performance with two approaches. However, we find a large improvement in the performance of the translate-test for swa and ewe with a boost of $12.4$ (GPT-4o comparison) and $10.4$ (LLaMa 3 70B comparison), respectively. For Swahili (swa), this improvement may be due to the strong machine translation performance on NLLB-200 , as reported in Adelani et al. [2] and Wang et al. [62].
AfriMGSM performance receives significant boost with translate-test evaluation
AfriMGSM is our most difficult benchmark in terms of overall performance on in-language evaluation. Table 4 shows that we can achieve a significant boost in performance with the translate-test on all languages we evaluated. We hypothesize that the current LLMs are better at reasoning in English than any other language. Interestingly, LLaMa 3 70B has the overall best performance boost of approximately $82\%$, and the performance boost confirms our hypothesis since it is more English-centric.
4.2 Few-shot results, Cross-Lingual Transfer, Sensitivity to Prompt Templates
Cross-lingual transfer in-language achieves better results
When there is large enough labeled data in English, we could leverage this cross-lingual signal for zero-shot evaluation. We trained on 400k English NLI examples, and we performed zero-shot transfer in in-language and translate-test setting. Table 5 shows that cross-lingual transfer using an Africa-centric smaller language model (AfroXLMR-76L) gave better results than prompting LLMs on average. AfroXLMR-76L has been pre-trained on all languages in IrokoBench, which explains the impressive performance. However, for multilingual encoders that have not seen some of the languages, prompting GPT-4o seems to be better, as shown in Appendix A.5.
Impact of few-shot vs zero-shot
Figure 2 shows the few-shot results for the IrokoBench datasets leveraging Aya-101, LLaMa 3 70B and GPT-4o when we provide few examples in in-language setting. We found out that Aya-101 LLM consistently benefited the least from additional few shots examples in all cases with only $+1.4$, $+2.0$ and $+0.1$ respectively. On the other hand, we see a large boost LLaMa 3 70B on both AfriXNLI and AfriMMLU tasks with $12.4$ and $6.5$ points improvement respectively. We hypothesize that additional examples in the language LLaMa 3 70B was not trained on were particularly helpful for the classification tasks. However, for mathematical reasoning tasks, this was unhelpful since LLaMa 3 70B is only able to reason properly in English as shown in Table 4. In general, GPT-4o benefits from adddtional few-shot examples, there is large improvement for AfriMGSM ($+10.5$) but smaller points improvement on AfriXNLI and AfriMMLU are smaller with $+2.0$ and $+0.4$ respectively.
One possibility for the difference in added performance from few-shot in part is due to the difference in context length. Aya-101 was by necessity built upon the mT5 [66] pre-trained base model given it was one of the few pre-trained models that had been trained on 101 languages. mT5 is relatively outdated given the rapid advances in LLM technology since its release in 2019. In particular, the context length of Aya-101 is only 1024 versus 8000 for LLaMa 3 70B and 32000 for GPT-4o. Hence, the truncation of some few-shot examples may disproportionately impact Aya-101.
Sensitivity to prompt templates
To understand whether sensitivity to prompts impacts results we perform an ablation for NLI where we evaluate results given two different prompts. Table 6 shows the results of two prompts we tested for: the first one is simpler [44], where the model is required to output true, false or neither, while the second prompt provides a detailed task description [37], and the output we are expecting is the label entailment. contradiction or neutral. We found out that open models like Aya-101 and Gemma 7B prefer the simpler prompt with Aya-101 achieving close to $50.0$ accuracy points on the task on average. However, closed models like Claude OPUS and GPT-4o prefer the Lai et al. [37] detailed prompt with $8.0$ and $7.0$ improvement in performance over the simpler prompt. In general, regardless of the prompt selection, GPT-4o is better than open models, but we can achieve better results for open models by providing a more detailed task description. While this evaluation is only for two prompts, an extensive prompt search would be important to explore all the tasks we evaluated.
(a) AfriXNLI
(b) AfriMMLU
(c) AfriMGSM
Figure 2: Few-shot evaluation on IrokoBench datasets, we performed 8-shots for AfriMGSM and 5-shots for the others.
5 Conclusion
In this paper, we introduced IrokoBench, a new benchmark for evaluating large language models (LLMs) on African languages. IrokoBench comprises three datasets focused on different tasks: natural language inference (AfriXNLI), multi-choice knowledge QA (AfriMMLU), and mathematical reasoning (AfriMGSM). Unlike previous benchmarks, which primarily involve simple text classification tasks, these datasets assessed the LLMs’ abilities in complex and knowledge-intensive areas. Our evaluation revealed a significant performance gap between high-resource languages (e.g., English and French) and African languages. Additionally, we observed a substantial disparity in performance between open models and proprietary models, with the latter generally outperforming the former, particularly in mathematical reasoning tasks. We hope that IrokoBench will serve as a valuable benchmark for evaluating future LLMs developed or adapted for African languages.
Limitations
Our benchmark has a few limitations: (1) The benchmark is human-translated which may include some translationese effects, it would have been better if they are all generated in the native African languages. However, this parallel translation allows us to evaluate and compare the same sentences in all these languages. (2) We only cover three language families in Africa, Nilo-Saharan, Austronesian, and Khoisan language groups are missing, one of the reason we excluded them is either lack of contact with professional translators or limited translation budget, we hope to extend to more languages in the future. (3) Prompts used for evaluation are not exhaustive some of the results may change with appropriate a more careful prompt search. The best prompt for different LLMs might be different.
Acknowledgment
This work was carried out with support from Lacuna Fund, an initiative co-founded by The Rockefeller Foundation, Google.org, and Canada’s International Development Research Centre. David Adelani acknowledges the support of DeepMind Academic Fellowship programme. This work was supported in part by compute credits from Oracle and Cohere For AI Research Grant. Also, we are grateful to OpenAI for providing API credits through their Researcher Access API programme to Masakhane for the evaluation of GPT-3.5 and GPT-4 LLMs. Finally, we are grateful to Hailey Schoelkopf and Lintang Sutawika for their assistance with the EleutherAI lm-eval tool.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g., for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3.5, more details in A.2
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 3.5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [N/A]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 3.2
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 3.2
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.3 for data and 3.4 for models
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 3.3
Appendix A Technical Appendix
Table 7: Languages covered in IrokoBench: including language family, region, number of L1 & L2 speakers, size of monolingual data on the web (in MADLAD corpus)
Language Family/branch Region # speakers # chars in MADLAD (MB) In Aya In BLOOMZ
English (eng) Indo-European / Germanic Across Africa 1457M 9,000, [^0] MB
French (fra) Indo-European /Romance Across Africa 310M 1,000, [^0] MB
Kiswahili (swa) Niger-Congo / Bantu East & Central Africa 71M-106M 2,400MB
Kinyarwanda (kin) Niger-Congo / Bantu East Africa 10M 749MB
Hausa (hau) Afro-Asiatic / Chadic West Africa 77M 630MB
Amharic (amh) Afro-Asiatic / Ethio-Semitic East Africa 57M 509MB
isiXhosa (xho) Niger-Congo / Bantu Southern Africa 19M 287MB
chiShona (sna) Niger-Congo / Bantu Southern Africa 11M 266MB
isiZulu (xho) Niger-Congo / Bantu Southern Africa 27M 257MB
Igbo (ibo) Niger-Congo / Volta-Niger West Africa 31M 251MB
Yorùbá (yor) Niger-Congo / Volta-Niger West Africa 46M 239MB
Sesotho (sot) Niger-Congo / Bantu Southern Africa 13M 227MB
Oromo (orm) Afro-Asiatic / Cushitic East Africa 37M 88MB
Luganda (lug) Niger-Congo / Bantu Central Africa 11M 48MB
Ewe (ewe) Niger-Congo / Kwa West Africa 7M 33MB
Twi (twi) Niger-Congo / Kwa West Africa 9M 25MB
Lingala (lin) Niger-Congo / Bantu Central Africa 40M 22MB
Wolof (wol) Niger-Congo / Senegambia West Africa 5M 5MB
A.1 Language covered
Table 7 provide the languages covered in the IrokoBench, their language family, regions located in Africa, number of speakers, and size of monolingual data available on the web based on the MADLAD cleaned corpus [36]—we only report number of characters in mega bytes. Additionally, we added an indication whether this language is covered in the pre-training of Aya-101 and BLOOMZ 7B LLMs.
A.2 Evaluation Tool
We use the EleutherAI LM Evaluation Harness (lm-eval) tool [15]—a popular evaluation tool that is helping to standardize LLM evaluation. The tool allows for three types of evaluation: log-likelihood, perplexity, and generation. The log-likelihood is more suitable for multiple-choice tasks since it helps to restrict the model’s option to fewer choices—more appropriate for weaker models. However, the log-likelihood approach cannot evaluate the generative capabilities of LLMs to generate coherent and relevant answers. Moreover, closed models are only accessible via API and do not provide access to the log probabilities, making it impossible to use the log-likelihood approach. To extract the correct answers for the task, we employed a verbalizer [25, 52]. For AfriMGSM, we used the default verbalizer provided by the tool. However, for AfriXNLI and AfriMMLU, we manually created a verbalizer for the closed models and used the log-likelihood request type for the open models.Table 8 provides the prompt templates used for evaluation present in the tool. For XNLI, we employed two types of prompts to understand the sensitivity of evaluation to type of prompt. We report on these results in Table 6.
Table 8: Prompt templates used for different tasks and datasets
. Task/Dataset Prompt AfriMGSM Question: {question} Answer: AfriMMLU You are a highly knowledgeable and intelligent artificial intelligence model answers multiple-choice questions about {subject} Question: {question} Choices: A: {choice1} B: {choice2} C: {choice3} D: {choice4} Answer: AfriXNLI (ANLI prompt [44]) {premise} Question: {hypothesis} True, False, or Neither? Answer: AfriXNLI (Lai’s prompt [37]) Please identify whether the premise entails or contradicts the hypothesis in the following premise and hypothesis. The answer should be exact entailment, contradiction, or neutral. Premise: {premise} Hypothesis: {hypothesis} Is it entailment, contradiction, or neutral?
(a) Amharic
(b) Yorùbá
(c) isiZulu
(d) Lingala
(e) Twi
(f) Wolof
Figure 3: Evaluation of AfriXNLI translations using AfriCOMET metric scores.
A.3 AfriCOMET metric scores for XNLI translation
We employ AfriCOMET evaluation metrics, as developed by Wang et al. [62], to automatically assess the quality of translations for our newly created benchmarks. Figure 3 depicts the histogram of scores obtained from AfriCOMET for AfriXNLI, illustrating promising results and offering compelling evidence for the effectiveness of our translations (Amharic, Yorùbá, isiZulu). However, the performance of this metric depends on if the language we are evaluating is covered in the pre-training of the base model of the metric i.e. AfroXLMR-large. In the case of Lingala, Twi and Wolof, the performance of the metric does not correlate with the human translation since they are not covered in AfroXLMR. Similar findings were reported in the original AfriCOMET QE paper that Twi had worse correlation with human judgement (i.e., 0.279 for Pearson, and 0.060 for Spearman) [62].
A.4 Task-specific results for all models
We provide the entire results of all LLMs on AfriMMLU, AfriXNLI and AfriMGSM tasks as shown in Table 9, Table 10 and Table 11
Table 9: AfriMMLU results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages. Best result in bold
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
FlanT5-XXL 48.4 37.6 22.4 22.6 24.2 27.0 24.8 26.6 22.6 27.4 25.5 25.2 28.4 27.2 26.2 24.8 25.6 25.8 25.4
mT0-XXL-MT 36.2 33.6 29.4 24.6 28.4 28.4 28.2 28.8 26.4 27.6 26.6 31.8 31.8 29.0 26.0 25.6 28.8 29.6 28.2
Aya-101 41.8 36.4 32.8 26.8 32.2 35.8 30.0 26.0 30.8 30.6 27.8 34.2 35.0 25.8 24.6 31.2 29.8 31.8 30.3
BLOOMZ 7B 33.0 29.2 20.6 25.0 22.2 23.6 23.2 24.2 23.0 25.4 24.8 23.0 26.0 19.5 23.4 23.2 25.6 24.0 23.5
Gemma 7B 51.2 47.0 26.6 27.2 29.0 31.0 28.2 31.0 30.4 30.8 29.4 29.2 34.0 28.4 28.8 27.0 29.0 32.4 29.5
LLaMa 2 7B 34.8 34.8 23.4 22.6 22.4 26.0 24.2 27.6 26.0 24.6 25.8 25.4 25.2 24.6 25.4 24.0 25.0 23.6 24.7
LLaMa 3 8B 49.0 45.6 26.8 25.8 29.6 29.4 28.4 28.8 22.4 29.2 29.4 27.0 30.4 24.8 27.6 24.6 28.0 28.0 27.5
LLaMa 3 70B 74.4 68.6 33.0 31.8 33.6 37.8 32.4 35.0 30.8 33.0 33.4 35.2 44.8 33.7 32.0 30.0 33.8 35.0 34.1
Command R 49.8 50.6 25.2 24.4 27.6 27.6 23.8 29.8 27.6 29.8 27.6 24.6 28.2 27.0 23.8 28.4 27.0 30.2 27.0
Command R+ 50.4 44.4 25.2 25.0 25.0 25.4 23.4 27.6 26.6 24.4 29.6 29.6 26.8 23.2 25.4 26.0 28.8 25.8 26.1
Claude OPUS 74.6 64.4 57.6 33.6 39.4 43.6 42.2 43.6 41.0 40.4 43.4 47.0 55.0 38.2 33.2 42.6 43.6 43.8 43.0
GPT-3.5 Turbo 61.8 56.4 30.6 29.0 33.4 32.6 32.0 31.8 29.2 31.0 34.4 32.4 40.6 29.4 29.2 29.4 30.8 33.0 31.8
GPT-4 Turbo 80.7 76.5 37.7 23.9 55.2 54.5 54.9 46.8 40.9 39.8 54.1 55.4 61.0 30.1 31.0 56.2 48.9 58.4 50.9
GPT-4o 78.6 71.4 55.2 31.6 57.4 55.8 56.8 55.4 45.0 48.8 58.4 62.0 67.4 37.6 32.6 60.8 50.8 60.6 52.3
Translate-Test (Eval. in English)
FlanT5-XXL 43.2 39.8 33.6 35.6 34.8 37.8 33.8 31.6 35.6 39.0 37.8 38.6 30.4 25.0 35.2 36.2 37.2 35.1
mT0-XXL-MT 34.0 31.0 27.2 30.6 28.0 27.4 27.6 25.8 28.2 31.2 31.2 29.0 30.4 26.6 29.4 30.2 28.0 28.9
Aya -101 37.2 32.0 30.4 30.8 31.6 29.2 32.6 29.2 29.0 33.4 33.8 31.6 30.8 25.4 32.8 32.6 32.4 31.1
BLOOMZ 7B 29.4 26.4 26.6 27.6 21.8 24.2 25.6 26.4 27.0 26.6 28.2 31.0 27.8 24.8 26.6 24.4 27.6 26.4
Gemma 7B 44.8 39.2 37.8 36.2 38.2 41.8 38.0 35.8 37.6 40.2 41.2 42.8 34.8 33.0 38.0 37.6 40.2 38.3
LLaMa 2 7B 33.0 31.0 29.8 29.0 32.6 29.4 29.6 27.4 30.0 27.4 30.2 31.4 28.6 28.6 34.2 28.4 30.6 29.9
LLaMa 3 8B 43.4 38.4 35.2 34.4 33.2 36.0 36.0 27.8 34.2 31.4 38.8 43.6 29.0 29.6 36.6 33.2 33.6 34.4
LLaMa 3 70B 64.4 54.4 42.2 44.4 48.6 47.2 49.8 41.2 49.6 48.4 54.0 60.0 41.4 29.8 49.6 49.2 49.2 47.4
Command R 47.4 41.6 31.8 36.2 36.6 36.8 38.2 29.0 37.8 34.0 43.0 44.8 31.4 28.8 42.0 35.0 36.8 36.5
Command R+ 44.0 36.4 31.8 35.4 37.4 36.6 35.8 34.2 35.6 37.6 37.0 44.0 31.6 26.6 37.4 37.2 36.6 35.7
Claude OPUS 63.2 52.4 39.8 46.2 48.8 49.2 46.2 40.2 45.6 48.0 52.4 75.3 37.6 31.0 51.2 47.4 50.2 47.6
GPT-3.5 Turbo 51.0 41.4 35.0 36.8 41.6 38.4 39.6 34.8 40.2 40.6 46.0 54.9 31.2 27.6 38.4 38.6 37.4 39.6
GPT-4 Turbo 61.0 50.0 37.6 43.2 47.6 48.2 46.8 40.4 44.6 46.2 53.0 72.7 36.4 28.0 48.4 46.6 49.0 47.0
GPT-4o 67.2 57.6 40.6 49.4 56.4 55.0 50.2 41.8 51.6 49.8 59.0 80.0 37.2 33.0 53.0 52.4 54.2 52.3
Table 10: AfriXNLI results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages. Best result per language in bold
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
AfroXLMR-76L 88.2 83.3 78.5 58.3 73.3 70.0 65.8 33.3 68.0 69.3 70.8 70.8 73.3 59.5 51.8 73.0 63.2 72.5 65.7
FlanT5-XXL 72.2 63.7 34.8 37.0 37.0 40.2 38.5 32.8 35.3 35.1 36.6 36.1 38.3 37.3 37.0 37.3 38.2 35.3 36.7
mT0-XXL-MT 63.5 61.5 58.3 38.5 57.5 56.5 51.3 33.5 54.2 47.2 54.8 56.0 55.5 48.8 39.7 56.3 52.3 55.0 51.0
Aya-101 61.5 60.5 58.2 42.2 56.8 54.7 48.7 34.5 44.5 51.7 55.8 54.7 55.2 48.2 37.3 53.7 48.3 54.3 49.9
BLOOMZ 7B 60.3 56.0 36.8 35.7 36.5 44.7 38.5 33.8 41.5 35.2 43.2 40.5 45.8 37.5 36.0 39.8 45.2 40.3 39.4
Gemma 7B 51.5 47.8 34.3 35.3 38.8 39.0 35.8 33.2 34.8 35.7 38.8 35.7 39.3 37.2 35.8 37.2 36.5 35.7 36.4
LLaMa 2 7B 50.0 48.5 33.7 33.3 33.2 33.5 33.5 33.8 33.5 33.3 32.8 33.5 34.0 34.2 31.2 34.0 33.3 33.3 33.4
LLaMa 3 8B 54.0 51.8 34.5 33.8 36.0 39.5 33.8 32.5 35.5 34.7 37.8 35.0 39.5 33.0 33.5 38.1 38.1 36.8 35.8
LLaMa 3 70B 51.3 50.2 36.8 33.3 35.3 36.5 34.5 33.0 33.5 33.3 33.8 34.3 36.5 33.0 33.5 34.0 33.0 33.0 34.2
CommandR 57.3 48.3 33.2 33.5 33.8 33.2 34.2 33.5 33.7 33.7 33.8 33.7 36.8 34.7 36.2 33.5 33.8 33.3 34.0
CommandR+ 56.5 51.5 41.0 33.5 38.8 40.5 38.2 33.0 37.8 34.0 37.0 37.7 43.0 36.2 36.5 37.2 43.2 39.5 37.9
Claude OPUS 85.7 74.7 61.3 54.5 61.5 45.0 64.5 30.5 63.7 50.2 57.0 68.3 70.5 56.0 50.0 68.8 63.7 63.7 58.1
GPT-3.5 Turbo 72.0 64.7 36.7 36.2 40.5 40.3 44.7 34.8 42.7 39.8 47.5 41.3 58.0 38.8 37.0 43.0 39.7 42.7 41.5
GPT-4 Turbo 86.0 82.2 65.2 38.7 69.3 69.5 67.7 32.2 65.0 58.3 67.5 65.7 67.2 47.5 39.7 72.0 65.3 69.2 60.0
GPT-4o 86.2 78.7 66.7 48.3 69.2 68.2 66.8 31.2 67.2 66.2 69.8 68.3 72.5 53.2 49.5 72.5 63.5 70.0 62.7
Translate-Test (Eval. in English)
AfroXLMR-76L 83.0 73.7 54.3 67.2 66.0 63.0 32.8 65.7 65.8 71.2 70.2 73.0 56.8 47.5 74.2 63.7 72.0 63.6
FlanT5-XXL 68.3 61.3 47.7 56.7 52.7 53.3 33.7 52.8 55.0 57.3 57.2 59.7 49.7 43.0 60.5 52.1 59.0 53.2
mT0-XXL-MT 59.3 54.5 45.3 52.7 50.0 49.8 34.5 48.2 50.2 55.0 53.0 56.8 46.2 42.7 54.8 50.7 54.3 49.9
Aya-101 60.3 55.2 47.0 51.8 52.2 50.3 33.0 49.0 49.0 52.0 53.3 54.0 45.5 41.5 54.5 48.8 54.0 49.4
BLOOMZ 7B 56.8 52.8 42.7 51.3 48.5 47.3 34.2 44.5 48.0 52.0 51.3 51.2 44.7 40.2 52.8 47.3 51.3 47.5
Gemma 7B 49.2 43.6 41.3 43.3 42.5 41.2 33.3 43.7 40.8 44.3 44.0 44.3 40.0 37.5 45.8 41.0 45.0 42.0
LLaMa 2 7B 48.0 46.7 39.3 43.7 43.0 42.2 34.0 42.0 41.8 42.3 43.7 45.7 42.0 39.2 46.3 40.5 44.3 42.3
LLaMa 3 8B 52.2 48.3 43.3 43.2 44.8 45.3 33.7 46.0 45.2 45.5 47.5 47.5 40.5 38.7 44.5 44.2 47.0 44.1
LLaMa 3 70B 48.5 42.8 41.3 43.7 39.8 40.3 33.3 43.3 39.7 41.1 43.0 44.3 39.7 38.7 42.3 38.2 42.7 40.9
CommandR 54.3 50.5 45.2 49.5 47.2 46.5 34.5 47.2 39.4 48.3 48.2 49.8 43.7 40.5 50.7 46.3 50.0 46.1
CommandR+ 55.0 51.2 42.5 47.7 45.3 45.3 35.2 46.2 43.8 48.0 47.8 49.5 41.3 39.5 47.3 46.2 49.7 45.4
Claude OPUS 56.0 65.7 45.7 76.3 59.0 55.0 32.0 57.7 55.8 61.5 62.5 63.8 46.3 41.8 63.5 55.3 59.8 56.4
GPT-3.5 Turbo 65.8 59.2 47.3 51.0 56.0 53.8 33.0 54.0 52.0 53.5 58.0 56.3 47.2 43.5 59.8 52.8 57.7 52.2
GPT-4 Turbo 75.2 64.0 43.0 52.8 59.0 55.8 31.7 51.0 52.5 55.8 58.5 60.2 45.0 39.7 63.2 52.5 59.5 52.8
GPT-4o 80.0 69.0 50.7 62.2 63.3 61.0 33.0 63.2 58.7 65.8 67.0 69.5 51.3 46.2 69.7 60.5 67.3 59.9
Table 11: AfriMGSM results in in-language and translate-test scenarios: Exact Match score per language. Average computed on only African languages. Best result per language in bold
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
FlanT5-XXL 24.0 22.0 7.6 2.8 4.4 4.0 5.6 2.0 4.4 2.4 2.4 2.8 10.4 1.2 2.4 1.6 3.2 3.6 3.8
mT0-XXL-MT 24.8 22.0 7.6 2.8 4.8 4.0 6.0 2.0 4.4 2.4 2.8 3.2 11.2 1.2 2.4 1.6 3.6 3.6 4.0
Aya-101 24.8 22.0 7.6 2.8 5.6 4.0 6.8 2.0 4.4 2.4 3.2 3.2 11.2 1.2 2.4 2.0 3.6 4.0 4.2
BLOOMZ 7B 24.8 22.0 7.6 2.8 5.6 4.0 6.4 2.0 4.4 2.4 2.8 2.8 11.2 1.2 2.4 2.0 3.2 4.4 4.1
Gemma 7B 24.8 22.0 7.6 3.2 4.4 4.4 5.6 2.0 4.4 2.4 2.4 3.6 11.2 1.2 2.4 1.6 3.2 4.0 4.0
LLaMa 2 7B 24.8 22.0 7.6 2.8 4.8 4.4 6.4 2.0 4.8 2.4 2.8 2.8 11.2 1.2 2.4 1.6 3.2 4.4 4.1
LLaMa 3 8B 24.8 22.0 7.6 2.8 4.8 4.4 5.6 2.0 4.8 2.4 2.4 3.2 11.2 1.2 2.4 1.6 3.2 3.6 4.0
LLaMa 3 70B 86.4 54.4 5.2 2.8 15.6 9.2 8.4 5.6 8.0 4.0 4.8 6.0 34.8 3.2 4.0 4.4 7.2 7.6 8.2
Command R 8.0 6.0 3.2 2.0 2.8 2.0 3.2 2.8 2.0 2.8 2.0 2.4 4.4 1.6 2.0 3.6 1.6 1.6 2.5
Command R+ 12.8 12.0 2.8 1.6 3.2 1.6 4.0 3.2 2.8 2.4 3.6 2.8 7.6 3.6 2.8 4.0 4.8 3.2 3.4
Claude OPUS 59.6 52.8 35.6 22.0 28.0 17.2 28.4 24.0 28.0 20.4 25.2 31.6 42.0 19.6 12.4 25.2 28.4 26.8 25.9
GPT-3.5 Turbo 60.0 43.2 1.2 0.8 5.2 1.2 5.6 2.8 7.2 3.2 2.8 3.6 35.2 3.6 2.8 3.2 2.0 2.4 5.2
GPT-4 Turbo 79.6 64.0 8.4 6.8 40.8 21.6 41.6 30.8 24.4 31.2 38.4 38.0 54.4 6.8 9.2 31.2 29.2 31.2 27.8
GPT-4o 58.4 43.6 4.4 5.6 42.0 32.0 36.4 30.8 22.8 39.2 33.2 28.8 44.8 10.4 12.4 29.2 30.8 28.4 27.0
Translate-Test (Eval. in English)
FlanT5-XXL 0.8 2.4 3.6 2.0 1.6 2.8 4.4 3.2 3.2 2.8 2.0 5.6 0.8 2.4 2.0 2.4 1.2 2.7
mT0-XXL-MT 3.6 2.4 1.6 1.6 1.6 3.2 2.4 2.8 2.4 2.4 2.4 3.6 1.6 2.8 4.0 2.8 4.0 2.6
Aya-101 8.0 8.0 6.0 7.6 4.0 7.2 5.2 4.4 5.2 8.4 8.0 10.8 4.4 5.6 6.0 8.4 6.4 6.6
BLOOMZ 7B 2.0 2.4 1.2 0.8 1.6 2.0 2.4 3.2 2.4 2.0 2.0 4.8 0.8 0.8 2.4 2.0 2.0 2.1
Gemma 7B 10.4 8.0 5.2 7.6 7.2 12.0 8.8 9.2 7.6 10.0 7.6 11.6 4.8 5.2 9.2 8.8 8.4 8.2
LLaMa 2 7B 9.6 4.0 5.2 3.6 7.2 7.2 6.0 6.4 6.0 7.2 6.0 9.2 5.2 3.2 6.8 6.4 5.2 5.9
LLaMa 3 8B 44.0 31.6 18.0 27.2 24.0 34.0 27.2 23.2 28.8 26.8 33.6 38.0 19.6 12.8 28.8 29.6 28.8 27.0
LLaMa 3 70B 69.6 53.6 27.6 48.4 40.4 54.0 47.6 34.0 47.6 46.8 54.0 69.6 29.2 20.8 50.0 49.2 51.2 45.3
Command R 9.6 8.0 2.4 4.8 4.8 7.6 4.8 5.2 6.8 5.6 8.4 8.8 5.2 2.8 4.0 6.0 7.2 5.8
Command R+ 11.2 7.2 5.6 9.6 8.8 10.4 10.8 8.0 9.6 7.2 10.4 13.2 7.2 5.6 10.4 10.0 12.4 9.2
Claude OPUS 47.2 38.8 23.2 33.2 22.0 41.2 36.0 26.4 32.4 36.0 40.8 48.4 19.6 16.8 35.6 36.4 36.4 32.7
GPT-3.5 Turbo 50.0 36.0 20.0 33.2 29.2 41.2 33.2 26.0 31.2 36.0 40.0 51.2 20.8 11.6 35.2 36.8 36.0 32.4
GPT-4 Turbo 63.6 42.0 20.4 39.2 35.2 47.6 41.6 27.2 38.8 39.2 46.0 56.8 23.2 12.4 41.2 42.0 40.4 37.1
GPT-4o 52.0 18.4 37.6 36.4 28.0 34.8 38.4 28.4 32.8 32.4 36.8 45.2 17.6 12.4 31.6 36.0 38.0 31.6
A.5 Cross-lingual transfer results for XNLI
In Table 12, we compare different multilingual masked language model (MLM) performance on African languages. XLM-R-large has 559M parameters and is trained on 100 languages, but only a few African languages are covered (amh, hau, orm, swa, and xho). Serengeti, on the other hand, has been pre-trained on all languages in IrokoBench, but it only has 240M parameters. AfroXLMR was adapted from XLM-R through continual pre-training on 17 African languages including 11 in IrokoBench (amh, hau, ibo, kin, orm, sna, sot, swa, xho, yor, and zul). AfroXLMR-76L follows the same technique by performing continual pre-training on XLM-R-large on 76 languages (72 African), all languages covered in IrokoBench are part of its pre-training.
We found Africa-centric MLM to perform better on average than massively multilingual models like XLM-R-large. Serengeti and AfroXLMR-base improved over larger-sized XLM-R-large by $+3.5$ and $+5.3$ points, respectively. Similarly, fine-tuning AfroXLMR-large, a larger version of AfroXLMR-base, results in an improved boost in performance with $11.8$ points. The best overall results were achieved by AfroXLMR-76L with a $16.1$ boost in performance over XLM-R-large. This is probably because all the languages are used in pre-training. We make use of AfroXLMR-76L has the baseline for all LLMs. Interestingly, we find GPT-4o to be competitive or better than other MLMs except the AfroXLMR-76L on average.
Table 12: Cross-lingual transfer results on AfriXNLI: We fine-tuned various multilingual encoders on English training data, and evaluated on other languages. Best result per language in bold
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
XLM-R-large 90.5 84.3 75.8 37.2 68.2 39.3 40.5 32.8 37.7 59.2 41.2 39.7 74.2 39.0 43.8 63.5 38.5 62.2 49.6
Serengeti 77.3 60.7 54.8 51.3 61.0 56.0 55.2 36.2 55.3 46.3 58.3 55.0 66.2 42.7 43.8 56.3 53.5 57.8 53.1
Afro-XLMR-base 81.5 78.5 71.3 36.3 68.8 59.8 57.7 37.8 44.2 56.7 59.8 61.5 67.0 40.8 41.0 61.0 52.2 62.7 54.9
Afro-XLMR-large 86.5 82.3 77.2 39.7 75.2 69.8 64.5 35.3 57.8 69.7 68.0 69.2 74.3 39.5 39.8 69.0 61.0 72.2 61.4
Afro-XLMR-76L-large 88.2 83.3 78.5 58.3 73.3 70.0 65.8 33.3 68.0 69.3 70.8 70.8 73.3 59.5 51.8 73.0 63.2 72.5 65.7
GPT-4o 86.2 78.7 66.7 48.3 69.2 68.2 66.8 31.2 67.2 66.2 69.8 68.3 72.5 53.2 49.5 72.5 63.5 70.0 62.7